A compound contains 12.8grams copper 6.4grams sulphur 12.8grams oxygen and 18grams of crystallisation find the simpless formulae of the compound


1
Expert's answer
2021-07-05T03:23:03-0400

Divide the above with their molar mass


For

"Copper = \frac{12.8}{63.546}=0.2"


"Sulphur = \frac{6.4}{32.065}=0.2"


"Oxygen = \frac{12.8}{15.999}=0.8"


"Crystallisation = \frac{18}{18.01528}=0.999"


Ratio

"\frac{0.2}{0.2}=1, \frac{0.2}{0.2}=1, \frac{0.8}{0.2}=4,
 \frac{1.0}{0.2}=5"


"CuS4O_25H_2O"